A He–Ne laser is made up of a cylindrical beam of light with a diameter of 0.750 cm. The energy is pulsed, lasting for 1.50 ns, and each burst contains an energy of 3.00 J.

(a) What is the length of each pulse of laser light? m

(b) What is the average energy per unit volume for each pulse? J/m3
Answer #1

a)

c = speed = 3 x 108 m/s

t = time = 1.50 ns = 1.50 x 10-9 sec

L = length

length is given as the distance travelled in a given time which can be written as

L = ct

L = (3 x 108) (1.50 x 10-9)

L = 0.45 m

b)

U = total energy carried = 3 J

Uavg = average energy per unit Volume = ?

d = diameter = 0.75 cm = 0.75 x 10-2 m

A = area = (0.25) \pid2 = (0.25) (3.14) (0.75 x 10-2)2 = 4.42 x 10-5 m2

average energy per unit Volume is given as

Uavg = U/(AL) = 3/((4.42 x 10-5) (0.45)) = 1.51 x 105 J